OpenStudy (anonymous):

Medals!!! Need more help...kindergarten children have normally distributed heights with a mean of 39 inches and a standard deviation of 2 inches. what is the probability that a randomly chosen kindergarten child will have a height between 37 and 41 inches?

OpenStudy (anonymous):

well, since they are both 1 deviation from the mean ...

OpenStudy (anonymous):

what does the empirical rule tell us?

ganeshie8 (ganeshie8):

more hints : 68-95-99.7

OpenStudy (anonymous):

about 68% lies within 1 SD

OpenStudy (anonymous):

of the mean

OpenStudy (anonymous):

+- 1 sd of the mean tha tis

OpenStudy (anonymous):

since 39 + 1sd = 41 and 39 - 1sd = 37 we are +-1 sd of the mean

OpenStudy (kirbykirby):

\[P\left( 37 < X < 41 \right)=\]\[P\left( \frac{37-39}{2} < \frac{X-39}{2} < \frac{41-39}{2} \right)=P\left( -1 < Z < 1 \right)=\]

OpenStudy (anonymous):

0.6826

OpenStudy (anonymous):

I really have no idea what I am doing

OpenStudy (anonymous):

do you know how to find the area underneath a curve?

OpenStudy (anonymous):

i see it in my book and try to work it out however i never end up getting the right answers. i have gone to almost every website.

OpenStudy (anonymous):

this is my last assignment for class, thankfully.

OpenStudy (anonymous):

the tables in a book are the calculations tha tpeople made a long time ago so that they would not have to do the calculus everytime they wanted to determine a value. instead of repeating the same thing over and over again, they just listed the values of greatest importance and looked it up when they needed it

OpenStudy (anonymous):

a normal curve is one that is high in the middle and skinny at the ends ... visually the total area under the curve is equal to 1, the area between any interval is equal to the probability of being in that interval